# Natural Language Processing RELIES on Linguistics

### v3

@@ 7 Study of Language
Our last facet highlights linguistics (and related areas) as the application domain. That is, those who study language, but are not necessarily computational linguists themselves, are a user base that can motivate NLP tasks and tools. [^11] Unlike the previous categories, which highlight ways that linguistics contributes to NLP, we view this category as bidirectional: language study motivates research towards developing NLP tools which can then contribute to the study of language. We focus here on NLP tools in the sense of off-the-shelf text analysis software, not NLP models and representations generally, but see e.g. Beguš et al. (2023); Mahowald et al. (2024); Chang and Bergen (2024) on the potential relevance of LLMs to broader linguistics and cognitive science. These language-system-focused applications can be distinguished from most contemporary NLP applications, which seek to make content more accessible through language (for example, by translating it, summarizing it, or reasoning about it).
@@
Classic cL tasks like parsing have applications to corpus linguistics: in order to study a linguistic pattern, it is often important to query not just strings, but with grammatical abstractions like tags, dependencies, and phrases. This was not the original goal of parsing researchers: parsing was a cL topic before the days of substantial digital corpora, and was studied to examine computational capacity to model natural language grammar. Now, the tools produced have found a secondary application in corpus linguistics, with parser outputs accessible through corpus search engines (Resnik and Elkiss, 2005; Ghodke and Bird, 2012; Krause and Zeldes, 2016; Guillaume, 2021; Hundt et al., 2012; Kulick et al., 2022). [^12] Search engines allowing users to query parsed corpora may also be motivated by information extraction applications (Shlain et al., 2020). The application to corpus linguistics perhaps motivates new angles on parsing research (e.g., syntactically searching a large text collection without having to pre-parse it in its entirety, or in a way that accounts for parser uncertainty).
@@
Documentary and historical linguistics also motivate and contribute to NLP advancement. In documentary and historical linguistics, data may be sparse, fragmented, primarily in image or audio form (without transcriptions), and lacking a standard orthography; the basic grammar of the language may still be a mystery; and the language may have few or no living speakers. (The case of endangered language documentation was discussed above in §3.) The exigencies of such settings call for considerably more noise-tolerance and interactivity than run-of-the-mill NLP tasks. Pure-text tasks like normalization (Robertson and Goldwater, 2018), as well as signal processing capabilities like OCR for historical documents (Berg-Kirkpatrick et al., 2013) or low-resource speech processing (Duong et al., 2016; Anastasopoulos et al., 2017; Liu et al., 2022), are important here. Still, algorithms and software infrastructure to assist humans engaging in multiple tasks collaboratively (elicitation, transcription, annotation, discovering grammatical generalizations, and developing usable resources for scholars and community members) in limited-data multimodal conditions are needed, and there is a long way to go before they can be used seamlessly by non-computational linguists (Gessler, 2022; Moeller and Arppe, 2024).
@@
Language teaching, as studied in the field of Applied Linguistics and practiced in classrooms as well as other modes of instruction and technological support, is another area ripe for further engagement with NLP. The research that exists in the NLP community tends to be formulated in narrow tasks like grammatical error correction (as a generation task) or essay scoring (as a regression task) (Burstein and Chodorow, 1999; Massung and Zhai, 2016; Wang et al., 2021; Klebanov and Madnani, 2022). Applications with richer contexts—for example, technologies that would monitor student progress over time and deliver adaptive pedagogical experiences—are more difficult to study, but could enhance student learning more holistically.
@@
Finally, we note the potential for NLP to embrace applications facilitating language-focused study in fields beyond linguistics proper, including literature Jakobson (1987), education, law Marmor (2014), communications Miller (1951), translation studies Gambier and Van Doorslaer (2009), history Piotrowski (2012), argumentation Cohen (1987), and lexicography McArthur (1998). In the NLP community, these fields are associated and developed with active workshops, e.g., LaTeCH-CLfL (CL for Cultural Heritage, Social Sciences, Humanities and Literature—8th iteration in 2024) (Bizzoni et al., 2024); BE(ducational)A (19th iteration) Kochmar et al. (2024); Argument Mining (11th iteration) Ajjour et al. (2024); and NL(egal)LP (3rd iteration) Aletras et al. (2024), among others. Important objectives featured in these workshops include creating structured resources, increasing interpretability, evaluating models and resources, and coping with low-resource settings, underpinning not only how language studies rely on NLP, but also how NLP relies on linguistics.
8 Challenges for the community
Having reviewed both obvious and subtle ways in which NLP relies on linguistic knowledge (highlighting Resources, Evaluation, Low-resource settings, Interpretability, Explainability, and Study of language), we now consider the future of NLP vis-à-vis linguistics.
A longstanding priority in NLP has been the development of linguistic resources including annotated datasets and linguistic processors (taggers, parsers, etc.). For a long time, automating analysis into linguistically-based representations was undertaken in part with the aim of paving the path towards NLU, building modules that capture features of language based on human conceptualizations (expressed in the form of manually crafted rules, or latent rules induced through training on annotations). Today, for some settings, LLMs offer an alternative path. However, there remain strong reasons for continuous improvement of linguistic resources. These reasons include some major engineering applications (Low-resource), but also goals that don’t immediately lie on the path towards NLU accuracy via scale. Among these other reasons are compute-constrained as well as targeted evaluation and data analysis settings, where taggers and parsers could shine with advantageous features such as efficiency, task-specificity (controllability) and precision. Moreover, for answering cL questions (as part of the study of language; section 7), we require accurate, fast, and robust linguistic resources.
Beyond continuing to develop cL resources, our review has highlighted other pertinent challenges for the NLP community, namely:
  • Efforts towards preserving the world’s languages. This requires cooperation between linguists, machine learning experts, and language communities (RLS).
  • Language learning applications and computational models of language acquisition, with potential for both the use and analysis of linguistic phenomena in NLP (ELIES).
  • Tasks that benefit from the integration of symbolic representations into models for interpretability and explanation (IE). This appears as a fruitful avenue for future work, while the broader application of neuro-symbolic modeling for performance is still an open debate (Hamilton et al., 2022; Shwartz, 2023).
  • Avenues where linguistics helps define and assess model facility and desiderata (Evaluation and Interpretability). Examples are linguistic calibration Mielke et al. (2022); Zhou et al. (2024), linguistic conventions Hawkins et al. (2020); Hua and Artzi (2024), and exploring metalinguistic behavior Behzad et al. (2023); Beguš et al. (2023); Hu and Levy (2023); Thrush et al. (2024). Research progress in these areas may also open doors to improving models.
  • As the quality of generated output improves, the focus of evaluation may shift increasingly away from linguistic form and toward function, including social meaning. We see great potential for sociolinguistics in creating resources and evaluation measures, studying the use of generated language in different social settings, understanding challenges, and helping develop solutions (Grieve et al., 2024) in all RELIES categories (and beyond).
9 Conclusion
@@
Given the recent emergence of powerful LLMs that leverage huge amounts of data to recognize patterns in text, the question of the relationship between natural language processing and linguistics, in particular how linguistic knowledge benefits work in natural language processing, is a pressing one. In this paper, we have illustrated the role of linguistics when compiling resources and conducting system evaluations (RE-); when building systems in low-resource settings (-L-); when pursuing granular interpretation or control of large-data systems (-IE-); and when connecting NLP to the study of human language (-S). While explicit linguistic knowledge is not necessary to achieve high accuracy on all NLP tasks, in this work we have highlighted the engineering utility of varying degrees of language knowledge.
We hope that this study will promote future work that leverages collaboration and connections between linguistics and computer scientists with the aim of NLP progress in diverse domains.
Acknowledgments
@@
Over the years, many scholars have influenced our perspectives on the relationship between NLP and Linguistics. We are grateful for feedback received in the course of writing this paper—from Julius Steen and Amir Zeldes on a draft, from Vivek Srikumar and Jena Hwang in a discussion, and Roland Kuhn on the problem of processing polysynthetic languages. We are also thankful for feedback that was received during invited presentations of this work at the 2024 NFDI plenary, and in Sowmya Vajjala’s NLP talk series.

### v4

@@ 7 Study of Language
Our last facet highlights linguistics (and related areas) as the application domain. That is, those who study language, but are not necessarily computational linguists themselves, are a user base that can motivate NLP tasks and tools. [^10] Unlike the previous categories, which highlight ways that linguistics contributes to NLP, we view this category as bidirectional: language study motivates research towards developing NLP tools which can then contribute to the study of language. To be clear, we regard the entirety of cL—spanning formal, descriptive, distributional/statistical, and experimental inquiries—as part of linguistics. (This includes the development and use of frameworks and resources of the sorts discussed in section 3.) These inquiries contribute to linguistic understanding by advancing accounts of linguistic diversity and the precision and robustness of theories. Naturally, NLP models and representations are part of the methodology of this research (in particular, the potential relevance of LLMs to broader linguistics and cognitive science has been widely debated: see Futrell and Mahowald (2025) for a review). For purposes of this section, though, we focus on specialized NLP software that can be applied to language studies, not NLP models and representations generally. These language-system-focused applications can be distinguished from most contemporary NLP applications, which seek to make content more accessible through language (for example, by translating it, summarizing it, or reasoning about it).
@@
Classic cL tasks like parsing have applications to corpus linguistics: in order to study a linguistic pattern, it is often important to query not just strings, but with grammatical abstractions like tags, dependencies, and phrases. This was not the original goal of parsing researchers: parsing was a cL topic before the days of substantial digital corpora, and was studied to examine computational capacity to model natural language grammar, highlighting issues like formal expressive power (Steedman, 2011). Now, the tools produced have found a secondary application in corpus linguistics, with parser outputs accessible through corpus search engines (Resnik and Elkiss, 2005; Ghodke and Bird, 2012; Bender et al., 2012; Hundt, Denison, and Schneider, 2012; Krause and Zeldes, 2016; Guillaume, 2021; Kulick, Ryant, and Santorini, 2022). [^11] Search in parsed corpora may also be motivated by information extraction applications (Shlain et al., 2020). The application to corpus linguistics motivates new angles on parsing research (e.g., syntactically searching a large text collection without having to pre-parse it in its entirety, or in a way that accounts for parser uncertainty).
@@
Documentary and historical linguistics also motivate and contribute to NLP advancement. In documentary and historical linguistics, data may be sparse, fragmented, primarily in image or audio form (without transcriptions), and lacking a standard orthography; the basic grammar of the language may still be a mystery; and the language may have few or no living speakers. (The case of endangered language documentation was discussed above in section 5.) The exigencies of such settings call for considerably more noise-tolerance and interactivity than run-of-the-mill NLP tasks. Pure-text tasks like normalization (Robertson and Goldwater, 2018), as well as signal processing capabilities like OCR for historical documents (Berg-Kirkpatrick, Durrett, and Klein, 2013) or low-resource speech processing (Duong et al., 2016; Anastasopoulos et al., 2017; Liu, Spence, and Prud’hommeaux, 2022), are important here. Algorithms and software infrastructure to assist humans engaging in multiple tasks collaboratively (elicitation, transcription, annotation, discovering grammatical generalizations, and developing usable resources for scholars and community members) in limited-data multimodal conditions are needed, and there is a long way to go before they can be used seamlessly by non-computational linguists (Gessler, 2022; Moeller and Arppe, 2024).
@@
Language teaching, as studied in the field of Applied Linguistics and practiced in classrooms as well as other modes of instruction and technological support, is another area ripe for further engagement with NLP. Much of the research in the NLP community has been formulated in narrow tasks like grammatical error correction (as a generation task) or essay scoring (as a regression task) (Burstein and Chodorow, 1999; Massung and Zhai, 2016; Wang et al., 2021; Klebanov and Madnani, 2022). Applications with richer contexts—for example, technologies that would monitor student progress over time and deliver adaptive pedagogical experiences—are perhaps more difficult to study, but could enhance student learning more holistically (Cui and Sachan, 2023; Qian et al., 2023; Glandorf et al., 2025; Bang et al., 2024).
@@
Finally, we note the potential for NLP to embrace applications facilitating language-focused study in fields beyond linguistics proper, including literature Jakobson (1987), education, law Marmor (2014), communications Miller (1951), translation studies Gambier and Van Doorslaer (2009), history Piotrowski (2012), argumentation Cohen (1987), and lexicography McArthur (1998). In the NLP community, these fields are associated and developed with active workshops, e.g., LaTeCH-CLfL (CL for Cultural Heritage, Social Sciences, Humanities and Literature—8th iteration in 2024) (Bizzoni et al., 2024); BEA (Building Educational Applications—19th iteration) Kochmar et al. (2024); Argument Mining (11th iteration) Ajjour et al. (2024); and NLLP (Natural Legal Language Processing–3rd iteration) Aletras et al. (2024), among others. Key objectives featured in these workshops include creating structured resources, increasing interpretability, evaluating models and resources, and coping with low-resource settings. This underscores the importance of interdisciplinary work between all of NLP, linguistics, and other fields from humanities and social sciences.
8 Conclusion and Outlook
@@
For a long time, automating analysis into linguistically-based representations—classic “cL“—was undertaken in part with the aim of paving the path towards NLU, building modules that capture features of language based on human conceptualizations (realized as manually crafted rules, or latent rules induced through training on annotations). Given the rise of LLMs that leverage huge amounts of data to recognize patterns in text, the question of the relationship between natural language processing and linguistics—in particular, how linguistic knowledge benefits work in natural language processing—is a pressing one. In this paper, we have illustrated the role of linguistics when compiling resources and conducting system evaluations (RE-); when building systems in low-resource settings (-L-); when pursuing granular interpretation or control of large-data systems (-IE-); and when connecting NLP to the study of human language (-S). While explicit linguistic knowledge is not necessary to achieve high accuracy on all NLP tasks, in this work we have highlighted the engineering utility of language knowledge.
Beyond continuing to develop cL resources and tools, our review has highlighted other pertinent challenges for the NLP community, including: Efforts towards preserving the world’s languages via cooperation between linguists, machine learning experts, and language communities (RLS); language learning applications and computational models of language acquisition, with potential for both the use and analysis of linguistic phenomena in NLP (ELIES); integration of symbolic representations and linguistic criteria generally into approaches that promote interpretability and explanation (IE).
Acknowledgements.
@@
Over the years, many scholars have influenced our perspectives on the relationship between NLP and Linguistics. We are grateful for feedback received in the course of writing this paper—from Julius Steen and Amir Zeldes on a draft, from Vivek Srikumar and Jena Hwang in a discussion, and Roland Kuhn on the problem of processing polysynthetic languages, as well as from reviewers. We are also thankful for feedback that was received during invited presentations of this work at the 2024 NFDI plenary, and in Sowmya Vajjala’s NLP talk series.
[…]